OpenStudy (anonymous):

−4(x + 1) − 3 = −3(x − 4)

OpenStudy (anonymous):

use distributive property on both sides: -4x - 4 - 3= -3x + 12 combine like terms on the left: -4x - 7 = -3x + 12 add 4x to both sides: -7 = x + 12 subtract 12 from both sides: -19 = x